# GraphTranslator: Aligning Graph Model to Large Language Model for Open-ended Tasks

### 11 Feb 2024 v1

@@ 3.2. Zero-shot Node Classification
The effectiveness is evaluated by utilizing several metrics: (1) Legality Rate, follow (Xuanhe Zhou, 2023), we use legality rate to measure the ratio that model produces valid answers. (2) Accuracy stands for the percentage of the correct predictions over total predicted samples. (3) Recall is calculated as the positive class recall rate in the 2 class task and macro-recall rate in the muli-class task. (4) F1-score is defined as the harmonic mean of precision and recall. The higher F1-score and legality rate hint the better effectiveness. (5) Top-k classification accuracy represents the percentage of the correct label is among the top-k predicted labels which is only adopted by the Arxiv sub-category prediction task. Note that the LLM-based models typically make predictions within the format of text rather than discrete labels, so we employ regular expression matching to extract the predicted class from response for evaluation. More details can be found in the Appendix D.
@@ 3.3. Graph Question Answering (GQA)
$\bullet$ Graph understanding: Taking the summarization of user’s interests in Q1 and friends’ interests in Q2 as examples, the correct responses require the model to recognize that both the user and friends are interested in cars. The Vanilla LLM  can understand the instructions, but struggles to extract the key characteristics. Specifically, in response A1, Vanilla LLM  lists numerous specific products as marked by the red strikethrough, which belong to friends, not users. And by A2, Vanilla LLM  summaries too many hobbies, but only car is the true shared interest. The reason is that the lengthy prompt in Q1 (over 1000 words) may disturb the attention mechanism of LLM. On the other hand, as highlighted in green, our GraphTranslator  identifies the primary interests and hobbies in A1, analyses the related personality and life needs, and in A2, successfully concludes that their common interest is cars. It demonstrates the superior ability of GraphTranslator  to extract and interpret graph information using the soft prompt translated from node embedding.
$\bullet$ Reasoning ability: The question Q3 requires models to understand graph information and reason why these people become friends. The Vanilla LLM  captures too many noise information of friends, thus tend to give a general explanation. Our GraphTranslator  provide explanations from three perspectives, i.e., the similarity of hobbies, the maintenance of friendship, and friend influence, which are all centered around the user and friends’ shared interest of cars. The final answer is fundamentally accurate, and it presents an explicit and logical reasoning process.
@@ Appendix A Discussion
$\bullet$ The Producer plays a pivotal role in dictating the overall quality of the Translator model. In future work, more topology information encoded in node embeddings, like node degree, can be included in the description in order to reduce information loss. Moreover, due to limited resources, we employed ChatGLM2-6B to generate descriptions in Producer. In future work, we can use larger-scale LLMs such as ChatGPT, to improve the quality of the generated text descriptions. And integrating novel LLM utilities and techniques, such as with Chain-of-Thought (Wei et al., 2022c) and AutoPrompt (Shin et al., 2020), also could further improve the performance.
@@ B.1. Graph Representation Learning
Graph representation learning, including earlier shallow graph embedding (Perozzi et al., 2014) and graph neural networks (Kipf and Welling, 2016; Velickovic et al., 2017; Hamilton et al., 2017), aims to capture and encode these relationships in a way that facilitates various downstream tasks. Traditional graph models primarily focus on supervised learning, requiring a substantial amount of labeled data for tasks such as node classification. These methods achieve strong performance when sufficient labeled data is available, but tend to underperform in scenarios with limited labeled data. Inspired by pre-trained language models, the new paradigm of ”Pre-train, Prompt, and Predict” has been recognized for its effectiveness in addressing few-shot downstream tasks  (Sun et al., 2022, 2023; Fang et al., 2022; Liu et al., 2023b; Huang et al., 2023; Zhou et al., 2023; Zhu et al., 2023). Despite their advancements, achieving zero-shot learning and open-ended tasks remains challenging.
@@ C.1. Code
  • LAVIS (https://github.com/salesforce/LAVIS): The logical architecture of LAVIS library served as the foundation for our code development.
  • ChatGLM (https://github.com/THUDM/ChatGLM-6B): An open-source LLM with the amazing language capabilities.
  • BLIP2 (https://arxiv.org/abs/2301.12597): our model is inspired from BLIP2.
In the spirit of open science and collaboration, we are also excited to share the source code of our work with the community. Our code repository will be published later.
C.2. Training Details.
C.3. Further Experiment
Appendix D Prompt Design
The prompts of Taobao and ArXiv datasets are presented in the Table 3, 4, 5, 6. For the prompts of the Taobao dataset, we translate them into English with Youdao  [^3] https://fanyi.youdao.com/. The prompts for the inference stage are shown in Table 5, we force the model to answer the number corresponding to the label in the Taobao dataset. Therefore, we employ regular expression matching to identify the relevant labels based on the digital number. For the ArXiv dataset, we extract labels based on the textual descriptions associated with each category.
@@ Appendix D Prompt Design
Table 6. Prompt Design For the GQA
Vanilla LLM
GraphTranslator
Question1
Please conduct some theoretical analysis of social impact based on the browsing, collecting, and purchasing behaviors of a user and his or her friends on an e-commerce platform. Here is some information about their behavior: {user_and_friends_behaviors} Please summarize the interests of this user.
¡Embedding¿. Please summarize the user’s interests
Question2
Please summarize the common interests and preferences of these friends
Please summarize the common interests and preferences of these friends
Question3
Why does this user become friends with these people?
Why does this user become friends with these people?
Evaluator(For ChatGPT)
The following is a 3-round dialogue. Please judge the quality of the responses based on the given criteria. The criteria are: A: The answers are correct and concise, the information is completely correct, and the reasoning is accurate. B: The answers are reasonable, with a very small number of information errors or imperfections. C: The answer is relevant to the question, but has obvious errors or inaccuracies in the content. D: The response is not relevant or completely invalid Here’s the conversation: [Round 1]Q:… A:… [Round 2]Q:… A:… [Round 3]Q:… A:… Please judge each answer one by one, output the rating results of 3 questions, and return in the form of a list. Format reference: [C, B, B]
The following is a 3-round dialogue. Please judge the quality of the responses based on the given criteria. The criteria are: A: The answers are correct and concise, the information is completely correct, and the reasoning is accurate. B: The answers are reasonable, with a very small number of information errors or imperfections. C: The answer is relevant to the question, but has obvious errors or inaccuracies in the content. D: The response is not relevant or completely invalid Here’s the conversation: Background: {user_and_friends_behaviors} [Round 1]Q:… A:… [Round 2]Q:… A:… [Round 3]Q:… A:… Please judge each answer one by one, output the rating results of 3 questions, and return in the form of a list. Format reference: [C, B, B]

### 28 Feb 2024 v4

@@ 3.2. Zero-shot Node Classification
The effectiveness is evaluated by utilizing several metrics: (1) Legality Rate, follow (Zhou et al., 2023b), we use legality rate to measure the ratio that model produces valid answers. (2) Accuracy stands for the percentage of the correct predictions over total predicted samples. (3) Recall is calculated as the positive class recall rate in the 2 class task and macro-recall rate in the muli-class task. (4) F1-score is defined as the harmonic mean of precision and recall. The higher F1-score and legality rate hint the better effectiveness. (5) Top-k classification accuracy represents the percentage of the correct labels which are among the top-k predicted labels (only for Arxiv dataset). Note that the LLM-based models typically make predictions within the format of text rather than discrete labels, so we employ regular expression matching to extract the predicted class from response for evaluation. More details can be found in the Appendix D.
@@ 3.3. Graph Question Answering (GQA)
$\bullet$ Graph understanding: Taking the summarization of user’s and friends’ interests (i.e., car) as examples, the Vanilla LLM  can understand the instructions, but struggles to extract the key characteristics. Specifically, in response A1, Vanilla LLM  lists numerous specific products as marked by the red strikethrough, which belong to friends, not users. And by A2, Vanilla LLM  summaries too many hobbies, but only car is the true shared interest. The reason is that the lengthy prompt in Q1 (over 1000 words) may disturb the attention mechanism of LLM. On the other hand, as highlighted in green, our GraphTranslator  identifies the primary interests and hobbies in A1, analyzes the personality and life needs, and in A2, successfully concludes that their common interest is cars. It demonstrates the superior ability of GraphTranslator  to extract and interpret graph information using the soft prompt translated from node embedding.
$\bullet$ Reasoning ability: The question Q3 requires models to understand graph information and reason why these people become friends. The Vanilla LLM  captures much noise information of friends, tending to give a general explanation. Our GraphTranslator  provides explanations from three perspectives, i.e., the similarity of hobbies, the maintenance of friendship, and friend influence, which are all centered around the user and friends’ shared interest of cars. The final answer is fundamentally accurate, and it presents an explicit and logical reasoning process.
@@ Appendix A Discussion
$\bullet$ The Producer plays a pivotal role in dictating the overall quality of the Translator model. In future work, more topology information encoded in node embeddings, like node degree, can be included in the description in order to reduce information loss. Moreover, due to limited resources, we employed ChatGLM2-6B to generate descriptions in Producer. In future work, we can use larger-scale LLMs such as ChatGPT, to improve the quality of the generated text descriptions. And integrating novel LLM utilities and techniques, such as with Chain-of-Thought (Wei et al., 2022b) and AutoPrompt (Shin et al., 2020), also could further improve the performance.
@@ B.1. Graph Representation Learning
Graph representation learning, including earlier shallow graph embedding (Perozzi et al., 2014) and graph neural networks (Kipf and Welling, 2016; Velickovic et al., 2017; Hamilton et al., 2017), aims to capture and encode these relationships in a way that facilitates various downstream tasks. Traditional graph models primarily focus on supervised learning, requiring a substantial amount of labeled data for tasks such as node classification. These methods achieve strong performance when sufficient labeled data is available, but tend to underperform in scenarios with limited labeled data. Inspired by pre-trained language models, the new paradigm of ”Pre-train, Prompt, and Predict” has been recognized for its effectiveness in addressing few-shot downstream tasks  (Sun et al., 2022, 2023; Fang et al., 2022; Liu et al., 2023b; Huang et al., 2023; Zhou et al., 2023a; Zhu et al., 2023). Despite their advancements, achieving zero-shot learning and open-ended tasks remains challenging.
@@ C.1. Code
$\bullet$ LAVIS (https://github.com/salesforce/LAVIS): The logical architecture of LAVIS library served as the foundation for our code development.
$\bullet$ ChatGLM (https://github.com/THUDM/ChatGLM-6B): An open-source LLM with the amazing language capabilities.
$\bullet$ BLIP2 (https://arxiv.org/abs/2301.12597): our model is inspired from BLIP2.
In the spirit of open science and collaboration, we are also excited to share the source code of our work with the community: https://github.com/alibaba/GraphTranslator.
@@
C.2. Training Details.
C.3. Further Experiment
Appendix D Prompt Design
The prompts of Taobao and ArXiv datasets are presented in the Table 3, 4, 5. For the prompts of the Taobao dataset, we translate them into English with Youdao  [^3] https://fanyi.youdao.com/. The prompts for the inference stage are shown in Table 5, we force the model to answer the number corresponding to the label in the Taobao dataset. Therefore, we employ regular expression matching to identify the relevant labels based on the digital number. For the ArXiv dataset, we extract labels based on the textual descriptions associated with each category.